OpenStudy (anonymous):

Solve the proportion t/16= 40/50 a. T=26 b. T=28 c. T=1.28 d. T=12.8

OpenStudy (goformit100):

(4/5) * 16

OpenStudy (anonymous):

t=(40/50)*16

OpenStudy (goformit100):

please do the further calculations

OpenStudy (anonymous):

t/16 = 40/50 cross multiply 50t = (40)(16) 50t = 640 t = 640/50 t = 12.8